OpenStudy (anonymous):

f(x)=9-4|x+1| Evaluate f(-5)

OpenStudy (photon336):

@phoebe_4567 what do you think should be the first step?

OpenStudy (mathmale):

Given the function f(x)=9-4|x+1|, please replace x with (-5) on both sides of your equation. Simplify. f(-5)=9-4|(-5)+1| = ??